Let L be the line given by <3,-1,2>+t<1,1,-1>,for tER.

1.show that the above line L lies on the plane -2x+3y-4z+1=0.

2.find an equation for the plane through the point p=(3,-2,4) that is perpendicular to the line <-8,2,0>+t<-3,2,-7>.


1
Expert's answer
2021-06-30T08:47:50-0400

1.


"x=3+t, y=-1+t, z=2-t"

"-2x+3y-4z+1"

"=-2(3+t)+3(-1+t)-4(2-t)+1"

"=-2(3+t)+3(-1+t)-4(2-t)+1"

"=5t-16"

Consider "t=0"


"5t-16=5(0)-16=-16\not=0"

Therefore the line L does not lie on the plane "-2x+3y-4z+1=0."


2.


"\vec n=\langle-3, 2, -7\rangle"

Point "P=(3,-2,4)."

The equation for the plane through the point "P=(3,-2,4)" that is perpendicular to the line "\langle-8, 2, 0\rangle+t\langle-3, 2, 7\rangle" is


"-3(x-3)+2(y-(-2))-7(z-4)=0"

or


"3x-2y+7z-41=0"